OpenStudy (anonymous):

need help solving -108=-18a

OpenStudy (anonymous):

Since u need to cancel out -18 to get a alone the opposite of multiplication is division so divide -18 out on each side of the equal sign

OpenStudy (anonymous):

All you need to do is divide -108 by -18, since the opposite of multiplication is division =)

OpenStudy (anonymous):

a=6

OpenStudy (anonymous):

Yes it does. Good job!

OpenStudy (anonymous):

thanks... 8-2=x+5

OpenStudy (anonymous):

Opposite of addition is subtraction

OpenStudy (anonymous):

am I subtracting by 5

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So the equation becomes 8-2-5 = x

OpenStudy (anonymous):

11x

OpenStudy (anonymous):

Not quite

OpenStudy (anonymous):

8-2 = 6 6-5 = ?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

how do I check answer

OpenStudy (anonymous):

You substitute x for the answer you got

OpenStudy (anonymous):

For example 8-2 = x+ 5 and you got x = 1 To check you answer you Plug in x 8-2 = 1 + 5

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

1/5x=-4/5

OpenStudy (anonymous):

need help with 1/5x=4/5

OpenStudy (anonymous):

Is the 1/5 next to the x or is it 1 over 5x

OpenStudy (anonymous):

next to the x

OpenStudy (anonymous):

In this case to cancel out the fraction u multiply by the reciprocal. Which is a fancy word for flipping the fraction. So 1/5 becomes 5/1 which is also 5

OpenStudy (anonymous):

So the equation becomes x = 4/5(5/1)

OpenStudy (anonymous):

do I have to multiply

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

20/5

OpenStudy (anonymous):

x=20/5

OpenStudy (anonymous):

Yes, now simplify the fraction

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thank you